# Understanding and Patching Compositional Reasoning in LLMs

### v1

@@ 1 Introduction
Compositional reasoning stands as a pivotal mechanism, unlocking the ability of learning systems to decompose complex tasks into manageable sub-tasks and tackle them step-by-step Lu et al. (2023); Lake and Baroni (2023). Despite the revolutionary impact of Large Language Models (LLMs) on the NLP landscape, they struggle at basic compositional reasoning tasks Dziri et al. (2023). This shortcoming is specifically highlighted by Press et al. (2023), who brought attention to the concerning “compositionality gap” in the realm of question-answering tasks. It was observed that there is a substantial failure rate of $\sim 40\%$ in two-hop compositional queries, even when they can successfully answer the individual single-hop queries that make up the two-hop question. Recent attempts improve the compositional reasoning capabilities of LLMs through carefully crafted prompting strategies developed by experts Wei et al. (2022); Zhou et al. (2023), enabling LLMs to autonomously rectify their compositional reasoning errors and continuously improve over time remains a largely unexplored frontier.
(a) Pre-Edit
(b) Post-Edit
Figure 1: Logit Lens inspecting results. x-axis refers to the layer; y-axis refers to the inspecting value (Eqn. 1). red and blue lines trace the implicit (association football) and explicit (England) reasoning results, respectively.
@@
This work, therefore, sets out to firstly delve into the specific failures to understand (RQ1) what accounts for these failures and (RQ2) which parts of the LLMs are responsible for them, and secondly develop strategies for patching these failures. Our initial step involves an analysis of a very recent dataset comprising compositional two-hop knowledge queries Zhong et al. (2023), selectively examining the cases where LLMs fail despite successfully answering the constituent single-hop queries. To ensure our findings and methodologies offer broad applicability, our analyses utilize two widely-used open-sourced LLMs: OpenAlpaca-3B Su et al. (2023b) and LLaMA-2-7B Touvron et al. (2023). Through meticulous examination of the failure instances, we identify three prevalent types of errors. Utilizing the Logit Lens tool nostalgebraist (2020), each error type highlights a critical shortfall in generating or leveraging the implicit reasoning result necessary for the explicit reasoning result [^1] Compositional two-hop queries require two-hop reasoning: implicit reasoning result is the first-hop reasoning result; explicit reasoning result is the second-hop reasoning result.. This gap is particularly concerning as it contrasts sharply with the intuitive two-hop reasoning process inherent to human cognition. An illustrative example of a “Hasty Answer” error is depicted in Figure 1(a), where the model prematurely concludes its reasoning without adequately incorporating the implicit reasoning result.
@@
Our main contributions and takeaways are summarized below: (1) successful compositional reasoning within LLMs hinges on its awareness of generating and leveraging implicit reasoning results; (2) MHSA modules in the middle layers (18/19-th layer) are significantly in charge of properly generating and leveraging implicit reasoning results; (3) by leveraging the second-hop computation graph as a reference for editing the located MHSA modules, CREME proves to be highly performing, on correctly answering not only the query used for editing $W_{o}^{l}$ but also the paraphrased queries and other compositional queries sharing the first-hop knowledge as well as maintaining little effect on irrelevant queries.
@@ 2 Background & Notation
Error type Input Implicit result Correct final result Predicted final result
@@
Incomplete Reasoning The head of state of the country where ORLAN holds citizenship is France Emmanuel Macron France
@@
Hasty Answer II The home country of the sport associated with Giorgio Chinaglia is association football England Italy
Table 1: Specific examples in $D_{gap}$ for three types of common errors. “Predicted final result” column refers to the wrong answers output by LLaMA-2-7B.
@@ 2.1 Logit Lens
where $L(h_{l}^{t},T_{tgt})$ measures how much information around $T_{tgt}$ is contained in $h_{l}^{t}$.
@@ 3 Analyzing Compositional Reasoning Errors
We query LLMs with all of compositional queries and the corresponding single-hop queries in $\mathcal{D}$. We filter out two subsets of $\mathcal{D}$: $\mathcal{D}_{single}$ and $\mathcal{D}_{gap}$. For each datum $(s_{1},r_{1},o_{1})\oplus(s_{2},r_{2},o_{2})$ in $\mathcal{D}$, $\mathcal{D}_{single}$ contains the datum where the both of $(s_{1},r_{1},o_{1})$ and $(s_{2},r_{2},o_{2})$ are successfully answered. Among $\mathcal{D}_{single}$, $\mathcal{D}_{gap}$ contains the datum where the answer for the compositional queries $(s_{1},r_{1},o_{1})\oplus(s_{2},r_{2},o_{2})$ are mis-predicted. [^2] Please find details in Appendix D.2. In our analysis of $\mathcal{D}_{gap}$, we have discerned a few common patterns shared among a substantial portion of the failures. Consequently, we have delineated three predominant types of errors, each characterized by distinct features, as outlined below. Distortion: LLMs fail to effectively generate implicit reasoning results in the reasoning process. The predicted answer for the first example in Table 1 is either United Kingdom or Italy. Considering both as countries (corresponding to nationality ($r_{2}$)), we conclude that the information about Donna Summer ($o_{1}$) distorts in middle hidden states. Incomplete Reasoning: LLMs directly output the first-hop reasoning result ($o_{1}$). In the second example of Table 1, LLaMA-2 outputs France ($o_{1}$) while the correct answer requires further reasoning. the head of state of ($r_{2}$) France ($o_{1}$) is Emmanuel Macron ($o_{2}$). Hasty Answer: LLMs predict the result without carefully reasoning. We further subdivide this type of errors into two categories: I: LLMs finally predict a close result based on the implicit reasoning result. For the third example in Table 1: LLMs predict Los Angeles or New York, both of which are famous city in the U.S.A., implying that LLMs manage to generate the implicit result ($o_{1}$:U.S.A.) while fails to incorporate “the capital of” ($r_{2}$) to generate final result $o_{2}$. II: LLMs take short-cut instead of step-by-step reasoning, leading to incorrect answers. Consider the fourth example in Table 1: the correct reasoning process should be (1): the sport associated with ($r_{1}$) Giorgio Chinaglia ($s_{1}$) is association football ($o_{1}$); followed by (2): the home country of ($r_{2}$) association football ($o_{1}$) is England ($o_{2}$). However, LLMs erroneously attribute Italy as the answer. This misstep is attributed to LLMs’ tendency to directly associate Giorgio Chinaglia ($s_{1}$) – noted for his Italian nationality – with the home country of the sport ($r_{2}$).
@@ 4 Analyzing the Inner Hidden States of LLMs for Compositional Reasoning
Figure 3: Logit Lens inspecting results with LLaMA-2-7B. (a) refers to the averaged result for inputs of compositional two-hop queries and (b) refers to the averaged result for second-hop queries. x-axis refers to the layer; y-axis refers to the 0-1 normalized probability. Yellow line and blue line refers to implicit results and explicit results respectively.
@@ 6.1 Dataset, Baseline and Evaluation Metric
The dataset $\mathcal{D}_{edit}$ we use for editing and evaluating LLMs is built based on the $\mathcal{D}_{gap}$ filtered in Section 3. For each example in $\mathcal{D}_{edit}$, it has the following fields: (1) Original input $I_{o}$ is a cloze test form of the compositional two-hop query. Accordingly, we also have the correct answer (ground-truth) and the originally predicted wrong answer for $I_{o}$: $A_{o}$ and $\widetilde{A_{o}}$, respectively666e.g.,for the fourth case in Table 1:$A_{o}$=England; $\widetilde{A_{o}}$=Italy.. In the experiment, we use $I_{o}$ and its corresponding second-hop query to edit the LLM. (2) Paraphrasing input $I_{p}$ is a paraphrase of $I_{o}$. Note that $A_{o}$ and $\widetilde{A_{o}}$ are also applicable to $I_{p}$. (3) Generalization input $I_{g}$ is a compositional two-hop query where its first-hop sub-knowledge is shared with $I_{o}$ while the second-hop sub-knowledge is different from $I_{o}$. We denote the correct answer for $I_{g}$ is $A_{g}$. (4) Irrelevant input $I_{i}$ is a compositional two-hop query that is irrelevant to $I_{o}$ and does not share the final answer with $I_{o}$. Detailed information about $\mathcal{D}_{edit}$ is available in Appendix B.
@@
In order to comprehensively validate the effectiveness of CREME, we propose four evaluation metrics: Correction, Paraphrasing, Generalization and Specificity. Following  Sakarvadia et al. (2023), all the metrics are formulated on the basis of Improvement Percentage (IP), which is calculated as $\text{IP}(I,A)=\frac{p_{\mathcal{M}^{*}}(A|I)-p_{\mathcal{M}}(A|I)}{p_{\mathcal{M}}(A|I)}$. This formula quantifies the enhancement in prediction probability of an answer $A$ given an input query $I$, facilitated by the post-edit LLM $\mathcal{M}^{*}$ in comparison to the pre-edit LLM $\mathcal{M}$. Specificially, Correction quantifies $\text{IP}(I_{o},A_{o})$ (larger is better); Paraphrasing is $\text{IP}(I_{p},A_{o})$ (larger is better); Generalization is $\text{IP}(I_{g},A_{g})$ (larger is better) and Specificity is $\text{IP}(I_{i},A_{o})$ (smaller is better). CoT-PatchScopes, due to its nature of input-dependent, only fits the Correction evaluation. We report the average results over $\mathcal{D}_{edit}$ in Section 6.2.
@@ 6.2 Experiment Results
Table 2: CREME versus baselines with the proposed four metrics: C for “Correction”, P for “Paraphrasing”, G for “Generalization” and S for “Specificity”.
@@
Method ROME (w. ground-truth) CREME (w.o. ground-truth)
Correction($\uparrow$) $\mathbf{98.0}\%$ $95.3\%$
Paraphrasing($\uparrow$) $62.5\%$ $\textbf{70.5}\%$
Generalization($\uparrow$) $+1.24$ $+\textbf{3.61}$
Specificity($\downarrow$) $+5.37$ $+\textbf{1.24}$
@@ 7 Related Work
LLMs fail to solve a large proportion of compositional multi-hop questions, even successfully solving all their single-hop sub-questions Press et al. (2023); Dziri et al. (2023). Early works towards mitigating this issue typically prepend crafted demonstration exemplars containing the “thought process” of solving the compositional query step-by-step and encourage LLMs to imitate the process via in-context learning Nye et al. (2021); Wei et al. (2022); Zhou et al. (2023); Drozdov et al. (2023); Press et al. (2023). Recent works turn to inspect the inherent compositional reasoning mechanism Hou et al. (2023) of LLMs.  Sakarvadia et al. (2023) manually injects implicit reasoning results into LLMs at the middle layers to correct compositional reasoning failures.  Ghandeharioun et al. (2024) fixes compositional reasoning errors through re-routing inner hidden representations in the computation graph to mimic chain-of-thought reasoning process. Nonetheless, their interventions in the reasoning process are rough so that the improvement is limited and hardly generalize to other related queries. To this end, we elaborately analyze the cause of compositional reasoning failures, locate a small set of parameters in LLMs that are responsible for such failures and precisely edit them to correct such failures.
@@ 8 Conclusion
We study the inner workings for the compositional reasoning of LLMs, which helps the black-box LLMs become more transparent and trustworthy Räuker et al. (2023). The CREME method introduced in this work is originally designed for correcting the compositional reasoning failures of LLMs. CREME only require slightly update a small set of parameters in LLMs and can generalize to a number of related queries (paraphrased queries or compositional queries sharing first-hop knowledge with the query used for conducting CREME). However, just like traditional model editing methods De Cao et al. (2021); Mitchell et al. (2022); Meng et al. (2022, 2023), it may also be utilized to insert inaccurate (or out-of-date) information into the pretrained LLMs.

### v2

@@ 1 Introduction
Compositional reasoning stands as a pivotal mechanism, unlocking the ability of learning systems to decompose complex tasks into manageable sub-tasks and tackle them step-by-step Lu et al. (2023); Lake and Baroni (2023). Despite the revolutionary impact of Large Language Models (LLMs) on the NLP landscape, they struggle at basic compositional reasoning tasks Dziri et al. (2023). This shortcoming is specifically highlighted by Press et al. (2023), who brought attention to the concerning “compositionality gap” in the realm of question-answering tasks. It was observed that there is a substantial failure rate of $\sim 40\%$ in two-hop compositional queries, even when they can successfully answer the individual single-hop queries that make up the two-hop question. Recent attempts improve the compositional reasoning capabilities of LLMs by making them decompose the question and explicitly output the step-by-step rationale with carefully crafted prompting strategies developed by experts Wei et al. (2022); Zhou et al. (2023). However, understanding the inherent mechanism of compositional multi-hop reasoning inside LLMs and enabling them to autonomously rectify their compositional reasoning errors and continuously improve over time remain largely under-explored.
Figure 1: An example of a typical compositional reasoning error pattern :Short-Cut(Section 3). Before patching, LLMs take short-cut reasoning by directly binding “The nationality” and “C. Auguste Dupin” (a fictional French detective) to incorrectly predict “France”. After patching, LLMs tend to firstly bind “the creator of” and “C. Auguste Dupin” to generate “Edgar Allen Poe” (implicit reasoning result) and then correctly predict “U.S.A.” (explicit reasoning result).
@@
This work, therefore, sets out to firstly delve into the specific failures to understand (RQ1) what accounts for these compositional reasoning failures and (RQ2) which parts of the LLMs are responsible for them, and secondly develop strategies for patching these failures. Our initial step involves an analysis of a very recent dataset comprising compositional two-hop knowledge queries Zhong et al. (2023), selectively examining the cases where LLMs fail despite successfully answering the constituent single-hop queries. To ensure our findings and methodologies offer broad applicability, our analyses utilize two widely-used open-sourced LLMs: OpenAlpaca-3B Su et al. (2023b) and LLaMA-2-7B Touvron et al. (2023). Through meticulous examination of the failure instances, we identify three prevalent types of errors. Utilizing the Logit Lens tool nostalgebraist (2020), each error type highlights a critical shortfall in generating or leveraging the implicit reasoning result necessary for the explicit reasoning result [^1] Compositional two-hop queries require two-hop reasoning: implicit reasoning result is the first-hop reasoning result; explicit reasoning result is the second-hop reasoning result.. This gap is particularly concerning as it contrasts sharply with the intuitive two-hop reasoning process inherent to human cognition. An illustrative example of compositional reasoning error is depicted in Figure 1, where the model incorrectly concludes its reasoning without properly generating and incorporating the implicit reasoning result.
@@
Our main contributions and takeaways are summarized below: (1) successful compositional reasoning within LLMs hinges on its awareness of generating and leveraging implicit reasoning results; (2) MHSA modules in the middle layers (18/19-th layer) are significantly in charge of properly generating and leveraging implicit reasoning results; (3) by leveraging the second-hop computation graph as a reference for editing the located MHSA modules, CREME proves to be highly performing, on correctly answering not only the query used for editing $W_{o}^{l}$ but also the paraphrased queries and other compositional queries sharing the first-hop knowledge as well as maintaining little effect on irrelevant queries [^2] Implementation will be available at https://github.com/Zhaoyi-Li21/creme..
@@ 2 Background & Notation
Error type Input Implicit result Correct final result Predicted final result Proportion
@@
Incomplete Reasoning The head of state of the country where ORLAN holds citizenship is France Emmanuel Macron France $36\%$
@@
Hasty Answer II The home country of the sport associated with Giorgio Chinaglia is association football England Italy $12\%$
Table 1: Specific examples in $D_{gap}$ for three types of common errors. “Predicted final result” column refers to the wrong answers output by LLaMA-2-7B. For each compositional knowledge, we query the language model with at most three paraphrased questions, and hence the predicted answers can be multiple. In the Proportion column, we calculate the proportions for each type of errors observed in the 200 two-hop examples from MQuAKE dataset Zhong et al. (2023).
@@ 2.1 Logit Lens
where $L(h_{l}^{t},T_{tgt})$ measures how much information around $T_{tgt}$ is contained in $h_{l}^{t}$. Note that typically there are multiple tokens in $T_{tgt}$ and we separately calculate the probabilities for all of these tokens and adopt their mean value to represent the whole $T_{tgt}$, similar with  Chanin et al. (2024). There are other works Sakarvadia et al. (2023); Yang et al. (2024) use the first token in $T_{tgt}$ to represent it, which sometimes can bring loss of information in $T_{tgt}$ and additional conflicts when processing targets like “United Kingdom” and “United States” as well.
@@ 3 Analyzing Compositional Reasoning Errors
We query LLMs with all of compositional queries and the corresponding single-hop queries in $\mathcal{D}$. We filter out two subsets of $\mathcal{D}$: $\mathcal{D}_{single}$ and $\mathcal{D}_{gap}$. For each datum $(s_{1},r_{1},o_{1})\oplus(s_{2},r_{2},o_{2})$ in $\mathcal{D}$, $\mathcal{D}_{single}$ contains the datum where the both of $(s_{1},r_{1},o_{1})$ and $(s_{2},r_{2},o_{2})$ are successfully answered. Among $\mathcal{D}_{single}$, $\mathcal{D}_{gap}$ contains the datum where the answer for the compositional queries $(s_{1},r_{1},o_{1})\oplus(s_{2},r_{2},o_{2})$ are mis-predicted. [^3] Please find details in Appendix D.2. In our analysis of $\mathcal{D}_{gap}$, we have discerned a few common patterns shared among a substantial portion of the failures. Consequently, we have delineated three predominant types of errors, each characterized by distinct features, as outlined below. Distortion: LLMs fail to effectively generate implicit reasoning results in the reasoning process. The predicted answer for the first example in Table 1 is either United Kingdom or Italy. Considering both as countries (corresponding to nationality ($r_{2}$)), we conclude that the information about Donna Summer ($o_{1}$) distorts in middle hidden states. Incomplete Reasoning: LLMs directly output the first-hop reasoning result ($o_{1}$). In the second example of Table 1, LLaMA-2 outputs France ($o_{1}$) while the correct answer requires further reasoning. the head of state of ($r_{2}$) France ($o_{1}$) is Emmanuel Macron ($o_{2}$). Hasty Answer: LLMs predict the result without carefully reasoning. We further subdivide this type of errors into two categories: I: LLMs finally predict a close result based on the implicit reasoning result. For the third example in Table 1: LLMs predict Los Angeles or New York, both of which are famous city in the U.S.A., implying that LLMs manage to generate the implicit result ($o_{1}$:U.S.A.) while fails to incorporate “the capital of” ($r_{2}$) to generate final result $o_{2}$. II: LLMs take short-cut instead of step-by-step reasoning, leading to incorrect answers. Consider the fourth example in Table 1: the correct reasoning process should be (1): the sport associated with ($r_{1}$) Giorgio Chinaglia ($s_{1}$) is association football ($o_{1}$); followed by (2): the home country of ($r_{2}$) association football ($o_{1}$) is England ($o_{2}$). However, LLMs erroneously attribute Italy as the answer. This misstep is attributed to LLMs’ tendency to directly associate Giorgio Chinaglia ($s_{1}$) – noted for his Italian nationality – with the home country of the sport ($r_{2}$). We calculate the proportions for each type of errors observed in the 200 two-hop examples of MQuAKE dataset Zhong et al. (2023) and post the statistics in Table 1 to demonstrate the commonness of these error types.
@@ 4 Analyzing the Inner Hidden States of LLMs for Compositional Reasoning
Figure 3: Logit Lens inspecting results with LLaMA-2-7B. (a) refers to the averaged result for inputs of compositional two-hop queries and (b) refers to the averaged result for second-hop queries. x-axis refers to the layer; y-axis refers averaged Logit Lens values after min-max normalization (i.e., the original values are linearly mapped to $[0,1]$). Yellow line and blue line refers to implicit results and explicit results respectively.
@@ 4.2 Verifying the Hypothesis via Intervention
(a) LLaMA-2-7B
(b) OpenAlpaca-3B
Figure 4: Intervention experiment: Brighter color indicates the intervention effect is more significant. In each subfigure, the upper row refers to the experiment group and the lower row refers to the comparison group. Note that for better visualization, we clip the effect value ($\leq 0$) to 00 for both of the experiment and comparison groups.
@@ 6 Patching Compositional Reasoning
What is worthy noting is that applying CREME only requires a single case of (compositional query, referenced second-hop query) while the patching effect can generalize to many other related cases. We showcase the effect of CREME in Table 2 and quantitatively discuss this generalization effect in Section 6.2.
Testing type Input Prediction w.o. CREME Prediction w. CREME
Hasty Answer II
Incomplete Reasoning
Table 2: Case study for generalization effect of correcting the (1) Hasty Answer II error: the original input (used for correcting) is “The country that the creator of C. Auguste Dupin belongs to is”. The original prediction is “France” (Reference: C. Auguste Dupin is French, while his creator Edgar Allan Poe. is American.); and (2) Incomplete Reasoning error: the original input (used for correcting) is “The capital of the country that Sven Väth is a citizen of is”. The original prediction is “Germany” (Reference: Berlin.).
@@ 6.1 Dataset, Baseline and Evaluation Metric
The dataset $\mathcal{D}_{edit}$ we use for editing and evaluating LLMs is built based on the $\mathcal{D}_{gap}$ filtered in Section 3. For each example in $\mathcal{D}_{edit}$, it has the following fields: (1) Original input $I_{o}$ is a cloze test form of the compositional two-hop query. Accordingly, we also have the correct answer (ground-truth) and the originally predicted wrong answer for $I_{o}$: $A_{o}$ and $\widetilde{A_{o}}$, respectively777e.g.,for the fourth case in Table 1:$A_{o}$=England; $\widetilde{A_{o}}$=Italy.. In the experiment, we use $I_{o}$ and its corresponding second-hop query to edit the LLM. (2) Paraphrasing input $I_{p}$ is a paraphrase of $I_{o}$. Note that $A_{o}$ and $\widetilde{A_{o}}$ are also applicable to $I_{p}$. (3) Generalization input $I_{g}$ is a compositional two-hop query where its first-hop sub-knowledge is shared with $I_{o}$ while the second-hop sub-knowledge is different from $I_{o}$. We denote the correct answer for $I_{g}$ is $A_{g}$. (4) Irrelevant input $I_{i}$ is a compositional two-hop query that is irrelevant to $I_{o}$ and does not share the final answer with $I_{o}$. Detailed information about $\mathcal{D}_{edit}$ is available in Appendix B.
@@
In order to comprehensively validate the effectiveness of CREME, we propose four evaluation metrics: Correction, Paraphrasing, Generalization and Specificity. Following  Sakarvadia et al. (2023), all the metrics are formulated on the basis of Improvement Percentage (IP), which is calculated as $\text{IP}(I,A)=\frac{p_{\mathcal{M}^{*}}(A|I)-p_{\mathcal{M}}(A|I)}{p_{\mathcal{M}}(A|I)}$. This formula quantifies the enhancement in prediction probability of an answer $A$ given an input query $I$, facilitated by the post-edit LLM $\mathcal{M}^{*}$ in comparison to the pre-edit LLM $\mathcal{M}$. Specificially, Correction quantifies $\text{IP}(I_{o},A_{o})$ (larger is better); Paraphrasing is $\text{IP}(I_{p},A_{o})$ (larger is better); Generalization is $\text{IP}(I_{g},A_{g})$ (larger is better) and Specificity is $\text{IP}(I_{i},A_{o})$ (smaller is better). CoT-PatchScopes, due to its nature of input-dependent, only fits the Correction evaluation. We report the average results over $\mathcal{D}_{edit}$ in Section 6.2. Note that to handling the multiple tokens in the answer, we calculate the probabilities for all of the tokens in the predicted probability distribution and report the mean value.
@@ 6.2 Experiment Results
To make readers have better sense of the realistic effect of using CREME to improve LLMs’ compositional reasoning performance, we also report the (1) decreasement percentage of $\log\text{PPL}$ values [^10] https://huggingface.co/docs/transformers/en/perplexity (of predicting correct answers) and (2) prediction accuracy in paraphrasing and generalization testing cases before and after applying CREME in Table 6.
@@
Table 3: CREME versus baselines with the proposed four metrics: C for “Correction”, P for “Paraphrasing”, G for “Generalization” and S for “Specificity”. Note that the values in the table is averaged improvement percentage (i.e., we calculate the improvement percentage for each single case and then do average over the dataset.) and hence it is meaningless to calculate values like $7.2\%\times(1+43.3)=318.96\%>100\%$.
@@ 7 Related Work
LLMs fail to solve a large proportion of compositional multi-hop questions, even successfully solving all their single-hop sub-questions Press et al. (2023); Dziri et al. (2023). Early works towards mitigating this issue typically prepend crafted demonstration exemplars containing the “thought process” of solving the compositional query step-by-step and encourage LLMs to imitate the process via in-context learning Nye et al. (2021); Wei et al. (2022); Zhou et al. (2023); Drozdov et al. (2023); Press et al. (2023). Recent works turn to inspect the inherent compositional reasoning mechanism Hou et al. (2023) of LLMs.  Sakarvadia et al. (2023) manually injects implicit reasoning results into LLMs at the middle layers to correct compositional reasoning failures.  Ghandeharioun et al. (2024) fixes compositional reasoning errors through re-routing inner hidden representations in the computation graph to mimic chain-of-thought reasoning process. Nonetheless, their interventions in the reasoning process are rough so that the improvement is limited and hardly generalize to other related queries. To this end, we elaborately analyze the cause of compositional reasoning failures, locate a small set of parameters in LLMs that are responsible for such failures and precisely edit them to correct such failures. The work and  Yang et al. (2024) are concurrent, where both of the two works show empirical evidence that LLMs can latently perform multi-hop reasoning internally.
@@ 8 Conclusion
We study the inner workings for the compositional reasoning of LLMs, which helps the black-box LLMs become more transparent and trustworthy Räuker et al. (2023). The CREME method introduced in this work is originally designed for correcting the compositional reasoning failures of LLMs. CREME only require slightly update a small set of parameters in LLMs and can generalize to a number of related queries (paraphrased queries or compositional queries sharing first-hop knowledge with the query used for conducting CREME). However, just like traditional model editing methods De Cao et al. (2021); Mitchell et al. (2022); Meng et al. (2022, 2023); Bi et al. (2024), it may also be utilized to insert inaccurate (or out-of-date) information into the pretrained LLMs, potentially resulting in negative influences in real-word applications of LLMs such as information retrieval Lian et al. (2020a) and recommendation Lian et al. (2014, 2020b).
Acknowledgements
Special thanks to Dr.Xiting Wang from Renmin University of China for her helpful comments to the early version of the work. We also sincerely thank the anonymous reviewers for their insightful suggestions to this work. The work was supported by grants from the National Key R&D Program of China (No.2021ZD0111801), the Research Grants Council of the Hong Kong SAR under Grant GRF 11217823 and Collaborative Research Fund C1042-23GF, the National Natural Science Foundation of China under Grant 62371411 and InnoHK initiative the Government of the HKSAR,Laboratory for AI-Powered Financial Technologies.